# Evaluating Quantumness, Efficiency and Cost of Quantum Random Number Generators via Photon Statistics

### v3

@@ VII Effect of $\mu,T,d$ on random number generation rate and cost
The entropy, which is the source of randomness of a QRNG, can be expressed as
$H(E)=H(Q)+H(C),$ (35)
where, $H(E),H(Q)$ and $H(C)$ are the total entropy generated by the RNG, the entropy of the quantum source and the entropy of the classical source respectively.
Definition 3.
An RNG is said to be perfect if $H(E)=H(U)$, where $H(U)$ is the entropy of uniform randomness. A QRNG is said to be perfect if $H(U)=H(E)=H(Q)$.
However, in the practical scenario, perfect QRNG is not possible. Therefore, the main goal in designing a QRNG is
@@
Figure 7: Entropy (38) of the probability distribution (29). When the product of the expected number of photons, $\mu T$, within an external reference cycle $T$, and the detection efficiency $d$ is very small, the entropy $H(Q)$ of the distribution generated by the QRNG is very close to the entropy $H(U)=8$ of the uniform random distribution. Therefore, if $\mu Td$ is very small, the random number generated by the QRNG would be close to the uniform random number.
The Shannon entropy Shannon (1948a, b) of a uniform distribution over $N$ points is given by
@@
A detailed derivation of this entropy has been added in Appendix B. Fig. 7 shows how the entropy (38) changes with $\mu Td$. Here we have chosen $N=256$. Therefore, the entropy (37) of the uniform distribution would be $H(U)=\log 256=8$. From Fig. 7, it can be easily seen that, if $\mu Td$ is very small, the entropy $H(Q)$ is very close to $8$, which is the entropy of the uniform distribution. In this section, by randomness we will refer to the randomness due to $H(Q)$ only.
VII.1 Effect of the Expected Photon Numbers
@@
VII.1.1 Effect of the Exected Photon Number on Cost
@@
VII.1.2 Effect of the Expected Photon Number on Random Number Generation Rate
A small value of $\mu$, which indicates a small number of photons, leads to a high chance of no photon detection in a particular external reference cycle of time interval $T$. That means those intervals will have no contribution to the final random number. Therefore, we require more cycles to produce random numbers of expected length. This will reduce the speed of the RNG. From Eq. (VI.2), the probability, $P(0)$, of no photon detection within time $T$ is given by
@@ VII.1.2 Effect of the Expected Photon Number on Random Number Generation Rate
VII.2 Effect of the timing error of the device on the External Reference Cycle
If we choose the external reference cycle $T$ very small, the product $\mu Td$ would also be small. This should increase the randomness, as discussed in Eq. (30) of the Subsection VI.2. However, a small $T$ would lead to a small length $\frac{T}{N}$ for the time bins. Now, if there are maximum $\delta_{t}$ errors in registering the photon-arrival-time, due to small time-bins, the time may be registered in a time-bin that is different from the actual time-bin as shown in Fig. 9. Let us consider that the actual detection happens in $i$-th bin, that is, within time interval $[(i-1)\frac{T}{N},i\frac{T}{N})$. If the actual detection happens within time interval $[(i-1)\frac{T}{N},(i-1)\frac{T}{N}+\delta_{t})$, the time may be registered as $(i-1)$-th bin. Similarly, if the actual detection happens within time interval $[i\frac{T}{N}-\delta_{t},i\frac{T}{N})$, the time may be registered as $(i+1)$-th bin. However, if the actual detection lies within the time interval $[(i-1)\frac{T}{N}+\delta_{t},i\frac{T}{N}-\delta_{t})$, there will be no error.
@@ VII.2 Effect of the timing error of the device on the External Reference Cycle
Now let us consider that $T$ is chosen to satisfy the relation
$\frac{T}{N}\geq k\delta_{t},$ (45)
for some suitable $k$. Then the probability, $p_{e}$, that the detection happens in $i$-th bin but outside of the interval $[(i-1)\frac{T}{N}+\delta_{t},i\frac{T}{N}-\delta_{t})$ is bounded above as $p_{e}\leq\frac{2}{k}$. Here we have assumed uniform distribution as the time $\frac{T}{Nk}$ is very small. Therefore, for a fixed error tolerance, $p_{tol}$, $k$ can be chosen as $k_{tol}=\frac{2}{p_{tol}}$. In that case, the minimum reference cycle $T_{min}$ should be
@@
VII.3 Effect of the Reference Cycle on Random Number Generation Rate and Cost
@@
VII.4 Effect of Detection Efficiency
@@
VII.4.1 Effect of the Detection Efficiency on Cost
@@
VII.4.2 Effect of the Detection Efficiency on Random Number Generation Rate
@@
VII.5 Performance and Cost of the QRNG
@@
Figure 11 is a surface plot of how the random number generation rate of the QRNG changes with $\mu$ and $d$. This shows that we have to consider high values of $\mu$ and $d$ to get a high rate. On the other hand, in Eq. (30) of Subsection VI.2 we show that to get maximum randomness from the quantum source, we have to minimize the product $\mu Td$ small. Therefore, this needs some optimization while choosing $\mu$ and $d$.
@@
VIII Results of Randomness Tests
As mentioned in Subsection II.3, the randomness of the random numbers is tested using statistical tests. Here we used four common statistical testing suites namely NIST, Dieharder, AIS-31 and ENT.
@@ VIII Results of Randomness Tests
The NIST Randomness Testing Suite is an extensive set of statistical tests developed by the National Institute of Standards and Technology (NIST) to evaluate the quality of randomness in random numbers. It is one of the most widely used suites for testing randomness. It contains $15$ different tests to examine different aspects of randomness. Most of the tests have the standard normal or $\chi^{2}$ as reference distribution for their test statistic. Here null hypothesis is taken as the number to be tested is random. For each test, the significance level is considered as $0.01$. If the value of the test statistic for some tests falls in the critical region, or equivalently, calculated $p$-value is less than $0.01$, then the corresponding tests fail rejecting the null hypothesis. A perfectly random number should pass all of the $15$ tests. As mentioned in Section V, the central limit theorem is applicable only for large $n$, tests that use the central limit theorem to determine the reference distribution, have a requirement for the bit stream, that is, the length of the testing sequence, to be large enough (at least $10^{6}$).
@@
IX Discussion
In this article, we consider two perspectives for analyzing QRNG. From the manufacturer’s point of view, we discuss how to determine when the source can be considered as quantum. Here we assume that photon detection time is available as a sample. We suggest two-fold statistical tests. The first one may be an interval estimation or hypothesis test to decide between sub-Poissonian and super-Poissonian. On the other hand, the last one, which is a $\chi^{2}$-test ensures Poisson distribution. This test can be performed for any quantum-photonic device to check its quantumness.
@@ IX Discussion
In practical cases, the dark count of a detector also contributes to the generation of random numbers. In our discussion, we did not consider the dark count. This work can be extended to see the effect of dark counts on the photon statistics. In our discussion, we have considered two models of photon arrival-time-based QRNGs. Similar analysis can be performed for other models of QRNGs.
@@ Acknowledgement
Appendix A Relationship between Non-classicality of Light, antibunching of Photons and sub-Poissonian Distributions
The second-order correlation function Fox (2006) of light is defined by
@@
where $n_{i}(t)$ is the number of counts registered on detector $i$ at time $t$. If the incoming light consists of a photon stream with long intervals between two consecutive photons, the product $n_{1}(t)n_{2}(t)$ is always 00, giving $g^{2}(0)=0<1$, which is not possible with classical lights. Thus, depending on $g^{2}(t)$, light is distinguished in the following three classes Fox (2006):
@@
The relation between photon count distribution and $g^{2}(\tau)$ is given by Zou and Mandel (1990)
@@
In the above equation, if $g^{2}(\tau)\leq 1$ for all $\tau$, it implies $\langle(\Delta n)^{2}\rangle<\langle N\rangle$, i.e., a sub-Poissonian statistics, indicating non-classicality of light. However, if $g^{2}(\tau)\leq 1$ for some $\tau$, there is no direct relation between the sub-Poissonian photon count statistic and the antibunching of photons. Xou and Mandel Zou and Mandel (1990) showed that if a source emits photons in two different nonvacuum modes, both having occupation number $n/2$, but different frequencies $\omega_{1}$ and $\omega_{2}$, the second-order correlation function takes the form
@@
where $\langle N\rangle$ and $\langle(\Delta N)^{2}\rangle$ are respectively the mean and the variance of the above distribution. Choosing $T=\frac{2\pi}{\omega_{1}-\omega_{2}}$, it can be seen that $\langle(\Delta N)^{2}\rangle<\langle N\rangle$, i.e., the light is sub-Poissonian although $g^{2}(0)=\frac{3}{2}-\frac{1}{n}>1$ for $n>2$ denoting the light is bunched. This example also shows that although both the antibunching of photons and the sub-Poissonian distribution of photon count indicate the non-classicality of light, they are not identical.
Appendix B Quantum Entropy for QRNG
The Shannon entropy Shannon (1948a, b) of the probability distribution given by Eq. (29) is given by
@@ Appendix B Quantum Entropy for QRNG
$=\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\left[\log\left(e^{\frac{\mu Td}{N}}\right)\frac{e^{-\frac{\mu Td}{N}}\left[1-(N+1)e^{-\mu Td}+Ne^{-(N+1)\frac{\mu Td}{N}}\right]}{(1-e^{-\frac{\mu Td}{N}})^{2}}-\log\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\right)\frac{e^{-\frac{\mu Td}{N}}(1-e^{-\mu Td})}{1-e^{-\frac{\mu Td}{N}}}\right]$
@@
This is the entropy generated by the photon arrival-time-based QRNGs we have discussed in this article.

### v4

@@ 7 Effect of $\mu,T,d$ on random number generation rate and cost
The randomness of an RNG is characterized by the entropy of the underlying probability distribution followed by the RNG. Let $H(Q)$ be the entropy generated by a quantum source. After postprocessing it becomes
$H(Q)\xrightarrow{postprocessing}H(E),$ (37)
where $H(E)$ is the final entropy generated by the QRNG. As deterministic classical post-processing cannot improve the entropy, in general $H(E)\leq H(Q)$.
Definition 2
An RNG is said to be ideal if $H(E)=H(U)$, where $H(U)$ is the entropy of uniform randomness. A QRNG is said to be ideal if $H(U)=H(E)=H(Q)$.
However, in the practical scenario, imperfections within used hardware components make ideal QRNG impossible. Therefore, the main goal in designing a QRNG is
@@
Figure 7: Total entropy (40) and minimum entropy (41) of the probability distribution (31). When the product of the expected number of photons, $\mu T$, within an external reference cycle $T$, and the detection efficiency $d$ is very small, $H(Q)$ and $H_{min}$ are very close to the entropy $H(U)=8$ of the uniform random distribution. Therefore, if $\mu Td$ is very small, the raw data generated by the QRNG would be close to the uniform random number.
The Shannon entropy [99, 100] of a uniform distribution over $n$ points is given by
@@
A detailed derivation of this entropy along with the definition of Shannon entropy and minimum entropy has been added in Appendix D. Clearly, $H(Q)\neq\log N=H(U)$. So, we have to apply post-processing as mentioned in Eq. (37) to achieve $H(E)=H(U)$ for some uniform distribution $U$ whose support contains less than $N$ points. This post-processing, also known as randomness extractor, is a data compression procedure using universal hashing [101]. The minimum entropy $H_{min}$ of the raw data decides the amount of this data compression. If $H_{min}\geq\frac{m}{n}$ for some positive integers $m,n$, an $n$-bit raw data needs to be compressed to an $m$-bit data to make it uniform random [17]. As $f(i)$ in Eq. (31) is a decreasing function, the minimum entropy of the photon arrival-time-based QRNG discussed above would be
$H_{min}$ $=-\log_{2}f(1)$
$=-\log_{2}\frac{1-e^{-\frac{\mu Td}{N}}}{1-e^{-\mu Td}}$
$=-\log_{2}\left(1-e^{-\frac{\mu Td}{N}}\right)+\log_{2}\left(1-e^{-\mu Td}\right).$ (41)
Nie et al. [17, Eq. (5)] gave a lower bound of this $H_{min}$. Since our aim is to reduce the amount of post-processing, which depends on $\mu Td$, we want to investigate $\mu,T$ and $d$ separately.
Fig. 7 shows how the total entropy (40) and minimum entropy (41) changes with $\mu Td$. Here we have chosen $N=256$. Therefore, the entropy (39) of the uniform distribution would be $H(U)=\log 256=8$. From Fig. 7, it can be easily seen that, if $\mu Td$ is very small, the entropy $H(Q)$ and $H_{min}$ are very close to $8$, which is the entropy of the uniform distribution. In this section, by randomness we will refer to the randomness due to $H_{min}$.
7.1 Effect of the Expected Photon Numbers
@@
7.1.1 Effect of the Exected Photon Number on Cost
@@
7.1.2 Effect of the Expected Photon Number on Random Number Generation Rate
A small value of $\mu$, which indicates a small number of photons, leads to a high chance of no photon detection in a particular external reference cycle of time interval $T$. That means those intervals will have no contribution to the final random number. Therefore, we require more cycles to produce random numbers of expected length. This will reduce the speed of the RNG. From Eq. (6.2), the probability, $P(0)$, of no photon detection within time $T$ is given by
@@ 7.1.2 Effect of the Expected Photon Number on Random Number Generation Rate
7.2 Effect of the timing error of the device on the External Reference Cycle
If we choose the external reference cycle $T$ very small, the product $\mu Td$ would also be small. This should increase the randomness, as discussed in Eq. (32) of the Subsection 6.2. However, a small $T$ would lead to a small length $\frac{T}{N}$ for the time bins. Now, if there are maximum $\delta_{t}$ errors in registering the photon-arrival-time, due to small time-bins, the time may be registered in a time-bin that is different from the actual time-bin as shown in Fig. 9. Let us consider that the actual detection happens in $i$-th bin, that is, within time interval $[(i-1)\frac{T}{N},i\frac{T}{N})$. If the actual detection happens within time interval $[(i-1)\frac{T}{N},(i-1)\frac{T}{N}+\delta_{t})$, the time may be registered as $(i-1)$-th bin. Similarly, if the actual detection happens within time interval $[i\frac{T}{N}-\delta_{t},i\frac{T}{N})$, the time may be registered as $(i+1)$-th bin. However, if the actual detection lies within the time interval $[(i-1)\frac{T}{N}+\delta_{t},i\frac{T}{N}-\delta_{t})$, there will be no error.
@@ 7.2 Effect of the timing error of the device on the External Reference Cycle
Now let us consider that $T$ is chosen to satisfy the relation $\frac{T}{N}\geq k\delta_{t}$, for some suitable $k$. Then the probability, $p_{e}$, that the detection happens in $i$-th bin but outside of the interval $[(i-1)\frac{T}{N}+\delta_{t},i\frac{T}{N}-\delta_{t})$ is bounded above as $p_{e}\leq\frac{2}{k}$. Here we have assumed uniform distribution as the time $\frac{T}{Nk}$ is very small. Therefore, for a fixed error tolerance, $p_{tol}$, $k$ can be chosen as $k_{tol}=\frac{2}{p_{tol}}$. In that case, the minimum reference cycle $T_{min}$ should be
@@
7.3 Effect of the Reference Cycle on Random Number Generation Rate and Cost
@@
7.4 Effect of Detection Efficiency
@@
7.4.1 Effect of the Detection Efficiency on Cost
@@
7.4.2 Effect of the Detection Efficiency on Random Number Generation Rate
@@
7.5 Performance and Cost of the QRNG
@@
Figure 11 is a surface plot of how the random number generation rate of the QRNG changes with $\mu$ and $d$. This shows that we have to consider high values of $\mu$ and $d$ to get a high rate. On the other hand, in Eq. (32) of Subsection 6.2 we show that to get maximum randomness from the quantum source, we have to minimize the product $\mu Td$ small. Therefore, this needs some optimization while choosing $\mu$ and $d$.
@@
8 Results of Randomness Tests
@@
Random numbers and their randomness testing have been discussed in Appendix E. The randomness of the random numbers is tested using statistical tests. Here we used four common statistical testing suites namely NIST, Dieharder, AIS-31 and ENT.
@@
The NIST Randomness Testing Suite is an extensive set of statistical tests developed by the National Institute of Standards and Technology (NIST) to evaluate the quality of randomness in random numbers. It is one of the most widely used suites for testing randomness. It contains $15$ different tests to examine different aspects of randomness. Most of the tests have the standard normal or $\chi^{2}$ as reference distribution for their test statistic. Here null hypothesis is taken as the number to be tested is random. For each test, the significance level is considered as $0.01$. If the value of the test statistic for some tests falls in the critical region, or equivalently, calculated $p$-value is less than $0.01$, then the corresponding tests fail rejecting the null hypothesis. A perfectly random number should pass all of the $15$ tests. As mentioned in Section 5, the central limit theorem is applicable only for large $n$, tests that use the central limit theorem to determine the reference distribution, have a requirement for the bit stream, that is, the length of the testing sequence, to be large enough (at least $10^{6}$).
@@
9 Discussion
In this paper, we consider two perspectives for analyzing QRNG. From the manufacturer’s point of view, we discuss how to determine when the source can be considered as quantum. Here we assume that photon detection time is available as a sample. We suggest two-fold statistical tests. The first one may be an interval estimation or hypothesis test to decide between sub-Poissonian and super-Poissonian. On the other hand, the last one, which is a $\chi^{2}$-test ensures Poisson distribution. This test can be performed for any quantum-photonic device to check its quantumness.
@@ 9 Discussion
In practical cases, the dark count of a detector also contributes to the generation of random numbers. In our discussion, we did not consider the dark count. This work can be extended to see the effect of dark counts on the photon statistics. Here, we have considered two models of photon arrival-time-based QRNGs, which fall in the category of trusted devices. A similar analysis can be performed for other QRNG models.
@@ Acknowledgement
Appendix A A Comparison between trusted QRNG and DI-QRNG
A device-independent quantum protocol uses quantum entanglement to ensure security. This can be verified using Bell inequality, quantum discord, etc. [71, 84, 85, 86]. On the other hand, if a protocol relies on single photons as a resource, the absence of entanglement prevents the application of the validation procedures mentioned above. However, we can still verify its quantumness using photon-count statistics [52, 53, 89, 90], discussed in Section  2 and Appendix B, to improve the security. A detailed comparison indicating advantages and disadvantages is given in Table 2.
Trusted QRNG Device-independent QRNG
Advantages of DI-QRNG 1. Required trusted assumptions. 1. Trusted assumptions are not required or partially required for semi-DI-QRNG.
2. Security relies on trusted assumptions. Hence, less secure. 2. Security relies on Bell violation or similar quantum principles [71, 84, 85, 86]. Hence, more secure.
Advantages of Trusted QRNG 1. No entanglement required. Hence it is cheaper. 1. Necessity of entanglement makes it costly.
2. No Bell test required. 2. Loophole-free Bell test is needed and performing a loophole-free Bell test is difficult [102, 103, 104, 105, 106].
3. Can be produced in the very compact form of USB or chip. 3. To avoid locality loophole in the Bell test, two entangled particles must be space-like separated [107, 108]. This forces the device to be either large or trusted during the Bell test.
4. Compatible for portable devices. 4. Incompatible in use cases with portable devices, e.g., automobiles, satellites, point of sale (POS) terminals, portable phones, laptops, etc. [36]
Table 2: A comparison between trusted and Device-independent QRNGs indicating their advantages and disadvantages.
From Table 2, one can easily conclude that although DI-QRNG provides more security, its size is not scalable to integrate it in a small space. Thus it is not suitable for many application scenarios requiring portable devices. Also, the high cost of DI-QRNG compared to trusted QRNG makes it incompatible with some uses like lottery games, token generations, etc.
Appendix B Relationship between Non-classicality of Light, antibunching of Photons and sub-Poissonian Distributions
The second-order correlation function [53] of light is defined by
@@ Appendix A A Comparison between trusted QRNG and DI-QRNG
where $n_{i}(t)$ is the number of counts registered on detector $i$ at time $t$. If the incoming light consists of a photon stream with long intervals between two consecutive photons, the product $n_{1}(t)n_{2}(t)$ is always 00, giving $g^{2}(0)=0<1$, which is not possible with classical lights. Thus, depending on $g^{2}(t)$, light is distinguished in the following three classes [53]:
@@
The relation between photon count distribution and $g^{2}(\tau)$ is given by [52]
@@
In the above equation, if $g^{2}(\tau)\leq 1$ for all $\tau$, it implies $\langle(\Delta n)^{2}\rangle<\langle N\rangle$, i.e., a sub-Poissonian statistics, indicating non-classicality of light. However, if $g^{2}(\tau)\leq 1$ for some $\tau$, there is no direct relation between the sub-Poissonian photon count statistic and the antibunching of photons. Xou and Mandel [52] showed that if a source emits photons in two different nonvacuum modes, both having occupation number $n/2$, but different frequencies $\omega_{1}$ and $\omega_{2}$, the second-order correlation function takes the form
@@
where $\langle N\rangle$ and $\langle(\Delta N)^{2}\rangle$ are respectively the mean and the variance of the above distribution. Choosing $T=\frac{2\pi}{\omega_{1}-\omega_{2}}$, it can be seen that $\langle(\Delta N)^{2}\rangle<\langle N\rangle$, i.e., the light is sub-Poissonian although $g^{2}(0)=(\frac{3}{2}-\frac{1}{n})>1$ for $n>2$ denoting the light is bunched. This example also shows that although both the antibunching of photons and the sub-Poissonian distribution of photon count indicate the non-classicality of light, they are not identical.
Appendix C Statistical Tools
C.1 Interval Estimation
C.2 Hypothesis Testing
In statistics, a hypothesis is a statement or an assumption about a population distribution. this may be right or wrong. Hypothesis testing is a procedure to decide whether a statistical hypothesis, called the null hypothesis $H_{0}$, should be rejected or not. A critical region or rejection region of hypothesis testing is a specific region of test statistic values, for which the null hypothesis is rejected. The test statistic for a hypothesis test is a function of the testing sample [109]. A hypothesis test is called one-tailed if the critical region appears only at one tail of the reference distribution of the test statistic. On the other hand, it is called a two-tailed test, if the critical regions appear at both tails of the above distribution.
A hypothesis test starts with two hypotheses: the null hypothesis ($H_{0}$) and the alternative hypothesis ($H_{a}$). Depending on the significance level $\alpha$, the critical region is decided. The acceptance or rejection decision of $H_{0}$ is based on what is called the $p$-value of the test statistic. It is defined as $P(|X|>|x|)$, where $X$ is a random variable, whose distribution is the reference distribution of the test statistic of a hypothesis test, and $x$ is the observed value of this statistic [110, 97]. When the observed test statistic falls within the critical region, $p$-value of the test statistic becomes less than $\alpha$. In this case, $H_{0}$ is rejected. If the $p$-value of the test statistic is greater than $\alpha$, the observed test statistic falls outside the critical region. In that case, $H_{0}$ will be accepted.
C.3 Goodness-of-fit Test
$H_{0}$: Sample data comes from a distribution to be tested.
$H_{a}$: Sample data not from that distribution.
Chi-square ($\chi^{2}$) test[111], Kolmogorov–Smirnov test (KS test) [112], and Anderson–Darling test [113] are some examples of the goodness-of-fit test [114]. Out of these tests, the last two tests are designed for continuous distributions only. However, the KS test is modified later to consider discrete distributions as well [115]. The $\chi^{2}$ test has been found more robust than the KS test for the continuous distributions [116].
Appendix D Quantum Entropy for QRNG
The entropy generated by an RNG is the source of randomness for its outcome. The outcomes of an RNG follow a probability distribution, say $P(X)$, where $X$ is the underlying random variable.
Definition 3
The Shannon entropy [99, 100] of a discrete random variable $X$ is defined as
$H(X)=-\sum_{x}P(X=x)\log P(X=x).$ (56)
The minimum entropy of an RNG is the smallest amount of the entropy that the RNG can produce.
Definition 4
Minimum entropy $H_{min}$ of raw data generated by a RNG is defined as
$H_{min}:=-\log_{2}\left(\max_{x}P(X=x)\right),$ (57)
where $X$ is the underlying discrete random variable.
This minimum entropy is used to decide the amount of post-processing required for transforming the raw data generated by an RNG to uniform random numbers.
The Shannon entropy of the probability distribution in Eq. (31) is given by
@@ Appendix D Quantum Entropy for QRNG
$=$ $\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\Bigg{[}\log\left(e^{\frac{\mu Td}{N}}\right)\frac{e^{-\frac{\mu Td}{N}}\left[1-(N+1)e^{-\mu Td}+Ne^{-(N+1)\frac{\mu Td}{N}}\right]}{(1-e^{-\frac{\mu Td}{N}})^{2}}$
$\;-\log\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\right)\frac{e^{-\frac{\mu Td}{N}}(1-e^{-\mu Td})}{1-e^{-\frac{\mu Td}{N}}}\Bigg{]}$
@@
This is the entropy generated by the photon arrival-time-based QRNGs we have discussed in this paper.
Appendix E Randomness Testing Suites
Definition 5
$X=x_{1}x_{2}\dots x_{n},x_{i}\in\{0,1\}\ \forall i,\text{ is true random}$
where $x=x_{j}x_{j+1}\dots x_{j+k}$.
To overcome this limitation, several randomness testing suites like NIST [65], Dieharder [66], AIS-$31$ [67], ENT [68] etc. have been proposed. Recently, Foreman et al. also suggested some test suite [117] for testing randomness. Each of these suits consists of multiple tests to decide the randomness of the input data. For each test, they calculate the value of the test statistic, and depending on these values the randomness is decided. NIST and Dieharder perform hypothesis testing with the null hypothesis that the number is random, AIS-31 performs interval estimation, and ENT directly returns the test statistic values. These suites consider different aspects, such as entropy, serial correlation, frequency of values, repetition of similar patterns etc. to perform the test. Some tests also perform Monte Carlo simulation [2] to validate the randomness.